What is the slope of the tangent to the curve y = x^2 – 6x + 9 at the point where x = 2
Answer by ewatrrr(24785)

 
Hi,
y = x^2 – 6x + 9
y ' = 2x+-+6 when x = 2 ⇒ y' = -2, the slope m of the tangent line
y = mx + b
y = -2x + b ||Using Pt(2,1) to find b |Note: when x = 2, y = 1
1 = -4 + b
5 = b
y = -2x + 5